Answer the following questions by referring to the equation below:

2 C6H10 + 17 O2  12 CO2 + 10 H2O

If 35 grams of C6H10 and 45 grams of oxygen are reacted,
i) which one is the limiting reagent?
ii) how many grams of carbon dioxide will be formed?
iii) how much of the excess reagent is left over after the reaction is complete?
iv) if 35 grams of carbon dioxide are formed, calculate the percent yield of this reaction.
1
Expert's answer
2019-10-01T03:23:40-0400

i) Number of moles of reagents must be estimated to determine the limiting reagent:

n = m/Mr,

where m - mass, Mr - molecular weight, n - number of moles.

From here:

n (C6H10) = 35 g / (82 g/mol × 2) = 0.2 mol

n (O2) = 45 g / 544 = 0.08 mol

Answer: O2 is a limiting reagent.

ii) As O2 is a limiting reagent, 0.08 mol of CO2 will be produced. From here:

m = n × Mr = 0.08 mol × 528 g/mol = 42.24 g.

Answer: 42.24 g of CO2.

iii) Only 0.08 mol of C6H10 will be used in the reaction. As a result, 0.2 mol - 0.08 mol = 0.12 mol of C6H10 remain after the reaction is complete. From here:

m = n × Mr = 0.12 mol × 164 g/mol = 19.68 g

Answer: 19.68 g of C6H10.

iv) The theoretical yield of CO2 is 42.24 g and actual yield is 35 g. The percent yield of the reaction is:

% yield = (actual yield / theoretical yield) × 100 % = (35 g / 42.24 g) × 100 % = 83 %.

Answer: 83 %.